What is the slope of a line perpendicular to the line whose equation is 12x−15y=−315

Answer by Cromlix(4381)

Sort 12x - 15y = -315 into y = mx + c
-15y = -12x - 315
15y = 12x + 315
y = 12/15x + 315/15
y = 4/5x + 21
When lines are perpendicular to one another
the gradients multiply together to give -1
m1 * m2 = -1
So, gradient of line perpendicular to
y = 4/5x + 21 = -5/4
Hope this helps.
:-)